A 1.00 L flask is filled with 1.15 g of argon at 25 ∘C. A sample of ethane vapor is added to the same flask until the total pressure is 1.250 atm .

Part A What is the partial pressure of argon, PAr, in the flask? Express your answer to three significant figures and include the appropriate units.

Part B What is the partial pressure of ethane, Pethane, in the flask? Express your answer to three significant figures and include the appropriate units.
Answer #1

The total pressure of flask

Ptotal = PAr + P ethane

the partial pressure of Ar also can becalculated from P=nRT/V

n (no of Moles of Ar) = wt of Ar / M.Wt of Ar = 1.15/40 = 0.0287

R = universal gas Constant = 0.082 L atm Mol-1 K-1 and T = 25 + 273 = 298K

V = Volume of Flask = 1 L

P = 0.0287 x 0.082 x 298 /1 = 0.7013 atm

P Ar = 0.7013 atm

The total pressure Ptotal = PAr + P ethane (P total = 1.250 atm)

P ethane = Ptotal - PAr = 1.250 - 0.7013 = 0.5486 atm

P ethane = 0.5486 atm